OpenStudy (anonymous):

The expression below is the factorization of what trinomial? (9x+8)(9x+3)

OpenStudy (anonymous):

okay medusa, lets take this one step at a time, but before we start, are you looking to learn this stuff or do you just want the answers?

OpenStudy (anonymous):

foil\[81x^2+27x+72x+24\]\[81x^2+99x+24\]

OpenStudy (anonymous):

9x*9x=81x^2 9x*3=27x 8*9x= 72x 8*3=24 than you add

OpenStudy (anonymous):

oh well ;)